The sum of the reciprocals of two consecutive positive integers is seventeen seventy-two. Write an equation that can be used to find the two integers. What are the integers?
Answer by nshah11(47)

Note that since the integers are positive and consecutive, let them be represented by x and (x + 1).
Reciprocal of x => 1/x
Reciprocal of (x + 1) => 1/(x + 1)
1/x + 1/(x + 1) = 1772
Getting a common denominator of x(x + 1):
(x + 1) + x = 1772(x)(x + 1)
2x + 1 = 1772x^2 + 1772x
1772x^2 + 1770x - 1 = 0
Via the quadratic formula:
x = (-1770 +/- √(1770^2 - 4(1772)(-1))/(2*1772) yields no integers at all. Pleas recheck your problem.